# Guobin Shen1,2,3∗ Dongcheng Zhao1,2,3,4∗ Linghao Feng3 Xiang He3Jihang Wang3Sicheng Shen3Haibo Tong3Yiting Dong3 Jindong Li3Xiang Zheng3Yi Zeng1,2,3,4†

### v1

@@ 1 Introduction
Large Language Models (LLMs), including architectures such as GPT brown2020language , Llama grattafiori2024llama , Qwen yang2024qwen2 , and Gemini deepmind_gemini , have achieved state-of-the-art performance across a wide range of natural language understanding and generation tasks. Their rapid deployment in real-world applications—from content creation and customer service to education and software development maeureka ; topsakal2023creating —highlights their transformative potential. However, as LLMs become increasingly embedded in safety-critical systems, ensuring their robustness and alignment has emerged as a paramount concern weidinger2022taxonomy ; wang2024navigating ; yao2024survey ; zhou2024multimodal ; zeng2024air ; dai2024safe .
A particularly acute threat to LLM safety is jailbreaking—a class of adversarial attacks in which carefully engineered prompts circumvent alignment constraints and elicit harmful, biased, or unethical outputs chu2024comprehensive ; zou2023universal ; liu2023jailbreaking ; yi2024jailbreak . Successful jailbreaks can trigger toxic language, misinformation, or even illegal instructions xie2024sorry ; yuan2024refuse ; shen2025jailbreak , undermining the guardrails built into state-of-the-art systems. Accordingly, the development of robust defenses and rigorous evaluation protocols for LLM jailbreak resistance has become an urgent research priority.
Despite valuable progress, current jailbreak evaluation approaches exhibit three key limitations. First, existing work often isolates individual components—such as attacks zou2023universal ; liu2024autodan ; chaojailbreaking or defenses xie2023defending ; wang2024defending ; shen2025jailbreak ; ji2024pku —without capturing their systemic interplay. Second, there is a lack of standardized benchmarking practices: evaluation protocols, datasets, and metrics remain fragmented zhang2025aisafetylab ; xu2024bag , which hinders reproducibility and fair comparison. Third, most evaluations are conducted on a limited scale, covering only a small subset of models, threats, or response evaluators zhang2024safetybench ; chao2024jailbreakbench . Moreover, critical factors such as computational cost, defense scalability, and the reliability of safety judges are often overlooked chang2024survey ; biarese2022advbench .
@@
JAILJUDGE liu2024jailjudgecomprehensivejailbreakjudge 2 3 18 4 4 (HF, OpenAI, Gemini, Claude)
EasyJailbreak zhou2024easyjailbreak 12 0 7 10 4 (HF, OpenAI, kimi, wenxinyiyan)
AISafetyLab zhang2025aisafetylab 13 16 7 1 3 (HF,vLLM, OpenAI)
HarmBench mazeika2024harmbench 18 0 3 33 5 (HF, vLLM, OpenAI, …)
@@
To address these challenges, we introduce PandaGuard, a unified and extensible evaluation framework that conceptualizes LLM jailbreak safety as a multi-agent system. In this formulation, attackers, defenders, target models, and safety judges interact within a structured ecosystem, as shown in Figure 1. PandaGuard abstracts and modularizes each component, supporting plug-and-play experimentation with 19 attack algorithms, 12 defense mechanisms, and multiple judgment strategies. This design facilitates controlled, reproducible evaluations and enables principled analysis of cross-component trade-offs in model safety. Built atop PandaGuard, we further develop PandaBench, a large-scale standardized benchmark suite encompassing approximately 3 billion tokens. PandaBench evaluates 49 open and closed-source LLMs—spanning various model sizes, release dates, and architectures—under diverse attack-defense combinations, as shown in Table 1. Beyond breadth, our framework supports practical deployment via flexible user interaction modes (attack, chat, serve) and compatibility with major inference engines including vLLM kwon2023efficient , SGLang zheng2024sglang , Ollama ollama_ollama , and other inference services accessible through standard APIs, thus enabling real-world usability and extensibility. Our contributions can be summarized as follows:
@@ 2 Background and Related Works
Jailbreak Attack Methodologies. Current attack methodologies can be categorized based on their access level and strategic approach. From an access perspective, white-box attacks leverage full knowledge of model parameters and architecture, utilizing gradient information to optimize adversarial prompts, with GCG zou2023universal pioneering gradient-based optimization of adversarial suffixes. In contrast, black-box attacks operate without access to model parameters, exemplified by PAIR chaojailbreaking and AutoDAN liu2024autodan . Strategically, jailbreak attacks have evolved from static templates liu2023jailbreaking ; jailbreakchat2025 ; wei2023jailbroken to more adaptive approaches, including proxy-based optimization methods li2024improved , zero-order alternatives like random search andriushchenko2024jailbreaking and genetic algorithms liu2024autodan , and semantic-level attacks that preserve malicious intent through linguistic transformations—such as PastTense andriushchenko2024does , Rainbow Teaming samvelyan2024rainbow , ArtPrompt jiang2024artprompt , and DeepInception li2024deepinception , which uses nested fictional characters to collectively work toward harmful goals, effectively bypassing safety mechanisms that focus primarily on token-level patterns.
@@
Defense Mechanisms. Defense mechanisms against jailbreak attacks span various implementation approaches and processing strategies. System-level defenses operate externally to the model, implementing input filtering jain2023baseline , response evaluation robey2023smoothllm ; wang2024defending , or in-context learning approaches xie2023defending ; zhang2024defending without requiring access to model parameters—including Self-Reminder xie2023defending , SmoothLLM robey2023smoothllm and its semantic variant ji2024defending , perplexity filtering jain2023baseline , paraphrasing techniques jain2023baseline , and SelfDefend phute2024llm . In contrast, model-level defenses directly modify the LLM’s parameters or internal representations, through approaches like representation engineering zou2023representation ; shen2025jailbreak , adversarial training mazeika2024harmbench ; ji2024pku , safety fine-tuning ji2024pku ; hsu2024safe , RLHF ouyang2022training , DPO rafailov2023direct , and Jailbreak Antidote shen2025jailbreak . An important but often overlooked aspect is the trade-off between security strength, computational overhead, and latency shen2025jailbreak ; liu2024autodan , which are critical for real-world deployment yet rarely evaluated systematically.
@@
Safety Evaluation. Evaluating LLM safety presents significant challenges, particularly in establishing consistent metrics and reliable judges. Current approaches include rule-based methods zou2023universal that often simply detect whether responses begin with refusals without necessarily assessing actual content harmfulness. LLM-based judges chaojailbreaking leverage other language models to evaluate response safety. Human evaluation remains the gold standard, though it is highly resource-intensive and difficult to scale. Recent studies have revealed concerning inconsistencies specifically with LLM-based judges wu2024meta ; gu2024survey , which can produce varying verdicts for identical inputs, raising questions about their reliability as safety arbiters. These stability issues underscore the need for more robust methodologies that can provide reliable assessments across diverse attack and defense scenarios.
@@
Existing Benchmarks and Limitations. Several benchmarks have emerged to standardize jailbreak evaluation, though each addresses only limited aspects of the safety ecosystem. JailbreakBench chao2024jailbreakbench provides a centralized repository of adversarial prompts, HarmBench mazeika2024harmbench implements various attacks and defenses, and SafetyBench zhang2024safetybench offers multiple-choice safety questions. EasyJailbreak zhou2024easyjailbreak evaluates the effectiveness of various attack methods against multiple models, and AISafetyLab zhang2025aisafetylab develops a tool for assessing model security when both attack and defense methods are employed. JAILJUDGE liu2024jailjudgecomprehensivejailbreakjudge establishes a jailbreak evaluation benchmark by integrating diverse attack scenarios and a multi-agent framework. Other contributions include PromptBench zhu2023promptbench , DecodingTrust wang2023decodingtrust , and TrustLLM sun2024trustllm , though they primarily evaluate static templates rather than adaptive attacks. Despite these valuable efforts, existing benchmarks suffer from key limitations: they often isolate specific attack vectors or defense mechanisms rather than examining their interplay, lack standardized algorithm implementations that lead to tight coupling between methods and models (obscuring true methodological contributions), conduct evaluations at insufficient scale, and overlook critical aspects such as computational overhead and judge reliability.
@@ 3 PandaGuard: A Framework for Safety-Critical LLM Evaluation
Component Abstraction and Implementation. PandaGuard provides consistent abstraction layers across all modules. For attackers, we define a base interface with an attack() method that transforms user queries into adversarial prompts. Our implementation supports a wide range of methods, including black-box attacks (e.g., PAIR chaojailbreaking , DeepInception li2024deepinception , AutoDAN liu2024autodan ) and optimization-based techniques such as GCG zou2023universal and RandomSearch andriushchenko2024jailbreaking .
Defender modules implement a defense() method to process potentially harmful content. We support three major paradigms: (1) detection-based methods (e.g., PerplexityFilter jain2023baseline ) that filter adversarial prompts, (2) prompt-based defenses (e.g., SelfReminder xie2023defending , GoalPriority zhang2024defending , SmoothLLM robey2023smoothllm ) that manipulate input phrasing, and (3) representation-level methods such as Jailbreak Antidote shen2025jailbreak that adjust internal model states to neutralize threats.
The target LLM interface supports both commercial API-based models (e.g., OpenAI, Anthropic, Gemini) and locally hosted models via frameworks like vLLM kwon2023efficient , SGLang zheng2024sglang , Ollama ollama , and Transformers wolf-etal-2020-transformers . Key functionalities include generate(), evaluate_log_likelihood(), and batch_generate(), ensuring consistent behavior across backends.
Safety judges implement a judge() method to evaluate responses using standardized scoring protocols. PandaGuard supports both rule-based judges zou2023universal and LLM-based judges chaojailbreaking ; mehrotra2024tree , enabling comparative analysis of judgment consistency and reliability.
@@ 4 PandaBench: Empirical Results and Key Insights
To comprehensively evaluate LLM jailbreak safety, we build PandaBench atop the PandaGuard framework. Unlike previous benchmarks that focus on limited models, isolated attack methods, or omit defense and judge considerations chao2024jailbreakbench ; biarese2022advbench , PandaBench offers the most comprehensive and reproducible jailbreak safety evaluation to date. PandaBench includes 49 diverse LLMs across model families and scales, 19 attack algorithms, 9 defense mechanisms, and multiple judging strategies. We adopt Attack Success Rate (ASR) as the primary metric, following the PAIR chaojailbreaking criterion—an attack is deemed successful only if the judge assigns a maximum score of 10.
To ensure fair comparison, we unify the proxy model used in attack and defense interactions. Specifically, we use Llama-3.1-8B grattafiori2024llama to generate adversarial prompts for attack algorithms and act as the agent for defense mechanisms. This eliminates discrepancies introduced by different backbone models and ensures that observed performance differences arise solely from algorithmic design. For a more comprehensive analysis, readers can refer to Appendix A for detailed experimental specifications, Appendix B for extended results with alternative judge implementations, and Appendix C for a discussion of limitations and future work. The full experimental configuration and results are available in our HuggingFace repository https://hf.co/datasets/Beijing-AISI/panda-bench.
@@ 4.2 Attack and Defense Mechanisms Analysis
Attack effectiveness across model landscape. Figure 3 reveals distinct patterns in attack method performance. RandomSearchRan andriushchenko2024jailbreaking consistently outperforms other techniques with an average ASR of 24%, followed by AIM jailbreakchat2025 (15%) and PastTense andriushchenko2024does (13%). This suggests that ensemble approaches combining multiple attack vectors often overcome diverse safety mechanisms more effectively than single-strategy attacks. We observe a clear effectiveness hierarchy among attack methods, with template-based approaches (AIM, BETTER_DAN) remaining surprisingly potent against many models, indicating persistent vulnerabilities in current alignment techniques. Proprietary models (Claude, GPT-4) demonstrate significantly higher resistance to jailbreaking compared to open-source alternatives, though certain attack-model combinations show unexpectedly high ASRs, suggesting specialized vulnerability patterns.
@@
Defense effectiveness across models. Figure 4 illustrates how different defense mechanisms perform across our model suite. All defenses reduce ASR compared to the Baseline, confirming their value in safety-critical deployments. Semantic-level defenses—Paraphrase jain2023baseline and Semantic SmoothLLM ji2024defending —demonstrate exceptional robustness by preserving meaning while neutralizing adversarial patterns. We observe that defense effectiveness varies substantially across models, with smaller open-source models benefiting more significantly from external defenses than larger proprietary ones, which have stronger inherent safety alignment. This suggests that defense strategies should be tailored to specific model characteristics rather than applied uniformly.
@@
Attack-defense interactions and practical trade-offs. Figure 5 examines the direct interactions between attack and defense methods and the associated implementation costs. The attack-defense heatmap (Figure 5a) reveals that semantic transformations maintain low ASRs even against sophisticated attacks by rewriting prompts while preserving intent. However, these strong defenses come with significant costs. Figure 5b shows that dialog-based defenses like SmoothLLM ji2024defending incur up to 5$\times$ higher token usage compared to the Baseline, raising practical deployment concerns. The green reference line represents the estimated token cost of post-generation safety filtering using PAIR chaojailbreaking ’s judge prompt, providing a cost-efficiency benchmark.
Beyond computational costs, Figure 5c demonstrates how defenses impact model utility using AlpacaEval alpaca_eval winrate. Stronger defenses tend to reduce performance more severely, with some methods degrading output quality by up to 25%. This reveals a critical three-way trade-off among safety effectiveness, computational efficiency, and task performance. For instance, SelfDefense phute2024llm excels in reducing ASR but causes notable performance drops, whereas PerplexityFilter jain2023baseline preserves high utility but offers weaker protection. Paraphrase jain2023baseline emerges as a practical middle ground with balanced trade-offs.
@@ 4.3 Safety Judge Reliability Analysis
Judge behavior variability. Figure 6a reveals substantial variations in how different judges evaluate the same model outputs. The rule-based judge (GCG zou2023universal ), which primarily detects refusal patterns, consistently reports significantly higher ASRs across all harm categories, defense methods, and attack strategies compared to LLM-based judges. This discrepancy underscores differing philosophies: whether jailbreaks are defined by refusal absence or by actual content harm.

### v3

@@ 1 Introduction
Large Language Models (LLMs), including architectures such as GPT [1], Llama [2], Qwen [3], and Gemini [4], have achieved state-of-the-art performance across a wide range of natural language understanding and generation tasks. Their rapid deployment in real-world applications—from content creation and customer service to education and software development [5, 6]—highlights their transformative potential. However, as LLMs become increasingly embedded in safety-critical systems, ensuring their robustness and alignment has emerged as a paramount concern [7, 8, 9, 10, 11, 12].
A particularly acute threat to LLM safety is jailbreaking—a class of adversarial attacks in which carefully engineered prompts circumvent alignment constraints and elicit harmful, biased, or unethical outputs [13, 14, 15, 16]. Successful jailbreaks can trigger toxic language, misinformation, or even illegal instructions [17, 18, 19], undermining the guardrails built into state-of-the-art systems. Accordingly, the development of robust defenses and rigorous evaluation protocols for LLM jailbreak resistance has become an urgent research priority.
Despite valuable progress, current jailbreak evaluation approaches exhibit three key limitations. First, existing work often isolates individual components—such as attacks [14, 20, 21] or defenses [22, 23, 19, 24]—without capturing their systemic interplay. Second, there is a lack of standardized benchmarking practices: evaluation protocols, datasets, and metrics remain fragmented [25, 26], which hinders reproducibility and fair comparison. Third, most evaluations are conducted on a limited scale, covering only a small subset of models, threats, or response evaluators [27, 28]. Moreover, critical factors such as computational cost, defense scalability, and the reliability of safety judges are often overlooked [29, 30].
@@
JAILJUDGE [31] 2 3 18 4 4 (HF, OpenAI, Gemini, Claude)
EasyJailbreak [32] 12 0 7 10 4 (HF, OpenAI, kimi, wenxinyiyan)
AISafetyLab [25] 13 16 7 1 3 (HF,vLLM, OpenAI)
HarmBench [33] 18 0 3 33 5 (HF, vLLM, OpenAI, …)
@@
To address these challenges, we introduce PandaGuard, a unified and extensible evaluation framework that conceptualizes LLM jailbreak safety as a multi-agent system. In this formulation, attackers, defenders, target models, and safety judges interact within a structured ecosystem, as shown in Figure 1. PandaGuard abstracts and modularizes each component, supporting plug-and-play experimentation with 19 attack algorithms, 12 defense mechanisms, and multiple judgment strategies. This design facilitates controlled, reproducible evaluations and enables principled analysis of cross-component trade-offs in model safety. Built atop PandaGuard, we further develop PandaBench, a large-scale standardized benchmark suite encompassing approximately 3 billion tokens. PandaBench evaluates 49 open and closed-source LLMs—spanning various model sizes, release dates, and architectures—under diverse attack-defense combinations, as shown in Table 1. Beyond breadth, our framework supports practical deployment via flexible user interaction modes (attack, chat, serve) and compatibility with major inference engines including vLLM [34], SGLang [35], Ollama [36], and other inference services accessible through standard APIs, thus enabling real-world usability and extensibility. Our contributions can be summarized as follows:
@@ 2 Background and Related Works
Jailbreak Attack Methodologies. Current attack methodologies can be categorized based on their access level and strategic approach. From an access perspective, white-box attacks leverage full knowledge of model parameters and architecture, utilizing gradient information to optimize adversarial prompts, with GCG [14] pioneering gradient-based optimization of adversarial suffixes. In contrast, black-box attacks operate without access to model parameters, exemplified by PAIR [21] and AutoDAN [20]. Strategically, jailbreak attacks have evolved from static templates [15, 37, 38] to more adaptive approaches, including proxy-based optimization methods [39], zero-order alternatives like random search [40] and genetic algorithms [20], and semantic-level attacks that preserve malicious intent through linguistic transformations—such as PastTense [41], Rainbow Teaming [42], ArtPrompt [43], and DeepInception [44], which uses nested fictional characters to collectively work toward harmful goals, effectively bypassing safety mechanisms that focus primarily on token-level patterns.
@@
Defense Mechanisms. Defense mechanisms against jailbreak attacks span various implementation approaches and processing strategies. System-level defenses operate externally to the model, implementing input filtering [45], response evaluation [46, 23], or in-context learning approaches [22, 47] without requiring access to model parameters—including Self-Reminder [22], SmoothLLM [46] and its semantic variant [48], perplexity filtering [45], paraphrasing techniques [45], and SelfDefend [49]. In contrast, model-level defenses directly modify the LLM’s parameters or internal representations, through approaches like representation engineering [50, 19], adversarial training [33, 24], safety fine-tuning [24, 51], RLHF [52], DPO [53], and Jailbreak Antidote [19]. An important but often overlooked aspect is the trade-off between security strength, computational overhead, and latency [19, 20], which are critical for real-world deployment yet rarely evaluated systematically.
@@
Safety Evaluation. Evaluating LLM safety presents significant challenges, particularly in establishing consistent metrics and reliable judges. Current approaches include rule-based methods [14] that often simply detect whether responses begin with refusals without necessarily assessing actual content harmfulness. LLM-based judges [21] leverage other language models to evaluate response safety. Human evaluation remains the gold standard, though it is highly resource-intensive and difficult to scale. Recent studies have revealed concerning inconsistencies specifically with LLM-based judges [54, 55], which can produce varying verdicts for identical inputs, raising questions about their reliability as safety arbiters. These stability issues underscore the need for more robust methodologies that can provide reliable assessments across diverse attack and defense scenarios.
@@
Existing Benchmarks and Limitations. Several benchmarks have emerged to standardize jailbreak evaluation, though each addresses only limited aspects of the safety ecosystem. JailbreakBench [28] provides a centralized repository of adversarial prompts, HarmBench [33] implements various attacks and defenses, and SafetyBench [27] offers multiple-choice safety questions. EasyJailbreak [32] evaluates the effectiveness of various attack methods against multiple models, and AISafetyLab [25] develops a tool for assessing model security when both attack and defense methods are employed. JAILJUDGE [31] establishes a jailbreak evaluation benchmark by integrating diverse attack scenarios and a multi-agent framework. Other contributions include PromptBench [56], DecodingTrust [57], and TrustLLM [58], though they primarily evaluate static templates rather than adaptive attacks. Despite these valuable efforts, existing benchmarks suffer from key limitations: they often isolate specific attack vectors or defense mechanisms rather than examining their interplay, lack standardized algorithm implementations that lead to tight coupling between methods and models (obscuring true methodological contributions), conduct evaluations at insufficient scale, and overlook critical aspects such as computational overhead and judge reliability.
@@ 3 PandaGuard: A Framework for Safety-Critical LLM Evaluation
Component Abstraction and Implementation. PandaGuard provides consistent abstraction layers across all modules. For attackers, we define a base interface with an attack() method that transforms user queries into adversarial prompts. Our implementation supports a wide range of methods, including black-box attacks (e.g., PAIR [21], DeepInception [44], AutoDAN [20]) and optimization-based techniques such as GCG [14] and RandomSearch [40].
Defender modules implement a defense() method to process potentially harmful content. We support three major paradigms: (1) detection-based methods (e.g., PerplexityFilter [45]) that filter adversarial prompts, (2) prompt-based defenses (e.g., SelfReminder [22], GoalPriority [47], SmoothLLM [46]) that manipulate input phrasing, and (3) representation-level methods such as Jailbreak Antidote [19] that adjust internal model states to neutralize threats.
The target LLM interface supports both commercial API-based models (e.g., OpenAI, Anthropic, Gemini) and locally hosted models via frameworks like vLLM [34], SGLang [35], Ollama [59], and Transformers [60]. Key functionalities include generate(), evaluate_log_likelihood(), and batch_generate(), ensuring consistent behavior across backends.
Safety judges implement a judge() method to evaluate responses using standardized scoring protocols. PandaGuard supports both rule-based judges [14] and LLM-based judges [21, 61], enabling comparative analysis of judgment consistency and reliability.
@@ 4 PandaBench: Empirical Results and Key Insights
To comprehensively evaluate LLM jailbreak safety, we build PandaBench atop the PandaGuard framework. Unlike previous benchmarks that focus on limited models, isolated attack methods, or omit defense and judge considerations [28, 30], PandaBench offers the most comprehensive and reproducible jailbreak safety evaluation to date. PandaBench includes 49 diverse LLMs across model families and scales, 19 attack algorithms, 9 defense mechanisms, and multiple judging strategies. We adopt Attack Success Rate (ASR) as the primary metric, following the PAIR [21] criterion—an attack is deemed successful only if the judge assigns a maximum score of 10.
To ensure fair comparison, we unify the proxy model used in attack and defense interactions. Specifically, we use Llama-3.1-8B [2] to generate adversarial prompts for attack algorithms and act as the agent for defense mechanisms. This eliminates discrepancies introduced by different backbone models and ensures that observed performance differences arise solely from algorithmic design. For a more comprehensive analysis, readers can refer to Appendix A for detailed experimental specifications, Appendix B for extended results with alternative judge implementations, and Appendix C for a discussion of limitations and future work. The full experimental configuration and results are available in our HuggingFace repository https://hf.co/datasets/Beijing-AISI/panda-bench.
@@ 4.2 Attack and Defense Mechanisms Analysis
Attack effectiveness across model landscape. Figure 3 reveals distinct patterns in attack method performance. RandomSearchRan [40] consistently outperforms other techniques with an average ASR of 24%, followed by AIM [37] (15%) and PastTense [41] (13%). This suggests that ensemble approaches combining multiple attack vectors often overcome diverse safety mechanisms more effectively than single-strategy attacks. We observe a clear effectiveness hierarchy among attack methods, with template-based approaches (AIM, BETTER_DAN) remaining surprisingly potent against many models, indicating persistent vulnerabilities in current alignment techniques. Proprietary models (Claude, GPT-4) demonstrate significantly higher resistance to jailbreaking compared to open-source alternatives, though certain attack-model combinations show unexpectedly high ASRs, suggesting specialized vulnerability patterns.
@@
Defense effectiveness across models. Figure 4 illustrates how different defense mechanisms perform across our model suite. All defenses reduce ASR compared to the Baseline, confirming their value in safety-critical deployments. Semantic-level defenses—Paraphrase [45] and Semantic SmoothLLM [48]—demonstrate exceptional robustness by preserving meaning while neutralizing adversarial patterns. We observe that defense effectiveness varies substantially across models, with smaller open-source models benefiting more significantly from external defenses than larger proprietary ones, which have stronger inherent safety alignment. This suggests that defense strategies should be tailored to specific model characteristics rather than applied uniformly.
@@
Attack-defense interactions and practical trade-offs. Figure 5 examines the direct interactions between attack and defense methods and the associated implementation costs. The attack-defense heatmap (Figure 5a) reveals that semantic transformations maintain low ASRs even against sophisticated attacks by rewriting prompts while preserving intent. However, these strong defenses come with significant costs. Figure 5b shows that dialog-based defenses like SmoothLLM [48] incur up to 5$\times$ higher token usage compared to the Baseline, raising practical deployment concerns. The green reference line represents the estimated token cost of post-generation safety filtering using PAIR [21]’s judge prompt, providing a cost-efficiency benchmark.
Beyond computational costs, Figure 5c demonstrates how defenses impact model utility using AlpacaEval [62] winrate. Stronger defenses tend to reduce performance more severely, with some methods degrading output quality by up to 25%. This reveals a critical three-way trade-off among safety effectiveness, computational efficiency, and task performance. For instance, SelfDefense [49] excels in reducing ASR but causes notable performance drops, whereas PerplexityFilter [45] preserves high utility but offers weaker protection. Paraphrase [45] emerges as a practical middle ground with balanced trade-offs.
@@ 4.3 Safety Judge Reliability Analysis
Judge behavior variability. Figure 6a reveals substantial variations in how different judges evaluate the same model outputs. The rule-based judge (GCG [14]), which primarily detects refusal patterns, consistently reports significantly higher ASRs across all harm categories, defense methods, and attack strategies compared to LLM-based judges. This discrepancy underscores differing philosophies: whether jailbreaks are defined by refusal absence or by actual content harm.